### v1

# Alpha Berkeley: A Scalable Framework for the Orchestration of Agentic Systems

Abstract
Coordinating workflows across heterogeneous control systems remains a central challenge in safety-critical environments such as scientific facilities, industrial plants, and energy infrastructures. Language-model–driven agents offer a natural interface for these tasks, but existing approaches often lack scalability, reliability, and human oversight. We introduce the Alpha Berkeley Framework, a production-ready architecture for scalable agentic systems that integrate conversational context with robust tool orchestration. The framework features dynamic capability classification to select only relevant tools per task, a plan-first orchestration model that generates execution plans with explicit dependencies and optional human approval, context-aware task extraction that combines dialogue history with external memory and domain resources, and production-ready execution environments with checkpointing, artifact management, and modular deployment. We demonstrate its versatility through two case studies: a tutorial-style wind farm monitoring example and a deployment at the Advanced Light Source particle accelerator. These results establish Alpha Berkeley as a reliable and transparent framework for agentic systems in high-stakes domains.
Figure 1: Overview of the Alpha Berkeley Framework workflow. Multi-turn conversational input and external data sources are first processed into a structured task. Relevant capabilities are dynamically classified on a turn-by-turn basis, and the selected tools are passed to the execution planner. The planner generates a complete, inspectable execution plan with explicit dependencies, which is then carried out by the agent with context tracking and artifact management.
I Introduction
Coordinating large-scale, safety-critical systems, such as smart electrical grids [1], industrial chemical processes [2], and scientific facilities such as particle accelerators [3], remains a central challenge in modern engineering. These environments integrate legacy and modern control systems that generate vast amounts of diverse data. Operators must act on this information under tight time constraints (for example, during accelerator fault recovery or power-grid disturbances) [4] yet retrieving, interpreting, and coordinating across systems remains cumbersome. Critical expertise is often preserved only as tacit human experience, scattered across institutional memory rather than codified in software. This tension between the need for flexibility and the rigidity of existing software architectures limits the ability to respond effectively to evolving operational demands and unexpected events.
Agentic AI [5], Language Model (LM) powered agents that orchestrate multi-step workflows through tool use, memory, and planning, offers a compelling opportunity to address these challenges. By providing a natural interface to data sources and control systems, LMs allow operators to express intent in high-level terms while delegating the orchestration of complex workflows to agents. In familiar domains such as web search [6] or travel planning [7], LMs already demonstrate this capability, as their training reflects widespread usage patterns and benchmark tasks. By contrast, in highly specialized domains such as particle accelerators, materials science beamlines, or other expert-driven facilities, much of the relevant expertise is absent from training corpora and cannot be inferred from general usage data. Instead, the knowledge is fragmented, highly context-dependent, and often confined to experts. Leveraging agentic AI in these environments therefore requires not only robust tool integration but also architectures capable of surfacing, formalizing, and operationalizing tacit human knowledge, while maintaining human-in-the-loop mechanisms to ensure safety and trust in high-stakes contexts.
In this paper, we introduce the Alpha Berkeley Framework [8], a production-ready architecture for building scalable agentic systems in complex scientific and industrial environments. Built on LangGraph [9] for its underlying graph-based orchestration substrate, our framework is distinguished by four characteristics:
  • Scalability through dynamic capability classification: On every interaction, the framework evaluates the current task and classifies available tools for relevance on a per-capability basis. This turn-by-turn filtering mitigates prompt explosion, prevents unnecessary tool loading, and allows the system to effectively generate a tailored “agent” for each prompt. As a result, the framework scales across large and growing tool inventories without overwhelming the language model.
  • Plan-first orchestration with human oversight: Unlike reactive agents that operate step-by-step, our framework generates complete execution plans upfront, with explicit input–output dependencies and error handling. Plans are presented to the user through integrated inspection and editing tools, enabling convenient approval and modification before execution begins.
  • Conversational awareness with external integration: The system incorporates multi-turn conversational context together with external resources, such as personal memory, databases, and APIs that are ensuring grounded, context-rich task extraction.
  • Production-ready adaptability: Designed for real deployments, the framework provides modular execution environments, artifact management, checkpointing, and diverse interfaces, including CLI, Open WebUI, and containerized services.
Together, these contributions move beyond proof-of-concept demonstrations toward a framework that is both flexible enough to adapt to new domains and robust enough for use in safety-critical operations. We illustrate the framework with two examples: a tutorial-style mock-up in a wind farm scenario, and a real-world deployment in the Advanced Light Source (ALS) control room. These case studies show both how the framework operates in practice and how it can be applied in complex, high-stakes environments.
II Background and Related Work
II.1 Agentic Systems and Tool Use
The foundations of agentic language models and multi-agent systems build on the classic definition of agents as entities that perceive their environment and act upon it, with rational agents aiming to maximize performance measures [10]. Extending this principle, multi-agent systems introduce shared environments in which agents may cooperate or compete. Early frameworks such as ReAct unified reasoning and acting in iterative loops for tool-based tasks [11], while AutoGen extended this paradigm toward multi-agent conversation orchestration, enabling specialized agents to collaborate on complex objectives [12].
Scaling tool use in LMs has since become a central concern. Toolformer demonstrated that LMs can be trained to make self-supervised API calls [13], while more recent work such as RAG-MCP addresses prompt bloat in tool selection by leveraging retrieval-augmented generation to select only relevant capabilities at inference time [14]. Beyond single-turn execution, memory and retrieval play an equally important role. Retrieval-Augmented Generation (RAG) integrates external knowledge into LM reasoning [15], but remains subject to context window limitations, such as the “Lost in the Middle” phenomenon [16]. To overcome this, memory-augmented agents like MemGPT [17] and subsequent surveys [18] have highlighted methods for extending long-context reasoning. Other work formalizes agent memory as a first-class capability, emphasizing the ability to retain past interactions to inform future steps [5]. Recent examples include SciBORG [19], a modular agentic framework for scientific tasks that employs FSA-based memory and context awareness to provide reliable execution and interpretable state transitions. Parallel efforts in task-oriented dialogue also offer complementary insights into grounded task extraction and intent modeling [20].
In summary, while prior work has advanced scaling of tool use and memory, these systems often still face challenges of prompt growth and domain grounding at inference time. We address these limitations in Section III.1 and Section III.2, where we introduce per-capability classification and grounded task extraction as core components of our framework.
II.2 Planning and Orchestration
Planning and orchestration in AI systems trace back to classical approaches such as STRIPS [21], PDDL [22], and hierarchical task network (HTN) planning [23]. Recent LM-based approaches build on these ideas by incorporating structured reasoning directly into language models, exemplified by Chain-of-Thought prompting [24] and Pre-Act, which combines multi-step planning with explicit reasoning steps [25]. DSPy further systematizes this direction, offering a modular framework for program synthesis and prompt optimization that reduces reliance on brittle, handcrafted prompt engineering [26, 27].
Complementing symbolic planning, graph-based architectures have emerged as powerful orchestration substrates for agents. LangGraph provides a task graph engine that enables explicit modeling of execution dependencies [9], while frameworks like PydanticAI contribute schema validation for structured outputs [28]. Related libraries, such as Instructor [29] and Outlines [30], offer constrained generation techniques that increase reliability in tool invocation.
Safety and human oversight remain critical in applying these systems to scientific and industrial domains. Reviews of self-driving laboratories emphasize that autonomy must integrate human intelligence for steering and safety [31], while approaches such as human-in-the-loop Bayesian experiment planning show how structured oversight can be embedded into experimental design [32]. More broadly, critical domains such as energy systems, particle accelerators, and laboratories require workflows with explicit approval points and safety boundaries.
Together, these observations motivate the plan-first architecture of our framework, detailed in Section III.3, where executable plans are generated with explicit dependencies and optional human review before execution.
II.3 Domain-Specific Applications
Agentic AI has seen rapid adoption in scientific and industrial applications, where domain-specific requirements demand specialized integrations. In the natural sciences, autonomous laboratories have been identified as a key driver of accelerated discovery [33], supported by robotic labs and LM-driven pipelines for materials exploration [34]. Within imaging sciences, the PEAR framework has demonstrated multi-agent LM automation for ptychography workflows [35], while other efforts have highlighted emergent autonomous research capabilities of LMs across chemistry and materials science [36]. Several domain-focused systems illustrate this diversity: ChemCrow augments LMs with chemistry-specific tools [37], Co-scientist enables autonomous experimental planning in chemistry [38], and CRISPR-GPT applies agentic orchestration to gene-editing workflows [39].
Among large-scale scientific facilities, agentic approaches have also been explored for particle accelerators and synchrotron beamlines. GAIA demonstrated an early accelerator operations assistant combining retrieval, scripting, and control [40], while VISION developed a modular AI assistant for beamline experiments, focusing on natural human–instrument interaction [41]. GAIA and VISION illustrate how agentic approaches can be applied to accelerators and beamlines, however, like many domain-specific systems, they rely on custom implementations that are difficult to scale. Our framework builds on these efforts with standardized capability integration, scalable orchestration, and production-ready reliability across both laboratory and facility contexts, as detailed in Section III.4.
III The Alpha Berkeley Framework
The Alpha Berkeley Framework [8] is designed to move beyond proof-of-concept agents and provide a scalable, production-ready architecture for domain-adaptable workflows. The conceptual sketch is shown in Figure 1. Its distinguishing features are a classification-based approach that keeps tool usage efficient even with large capability inventories, upfront execution planning with optional human oversight, and conversational awareness that integrates multi-turn context with external memory and databases. At the same time, the framework is engineered for deployment rather than demonstration: it offers modular execution services, checkpointing and recovery for reliability, and multiple interface options ranging from a command line to browser-based GUIs and containerized services. Together, these elements make the framework both adaptable to new use cases and robust enough for sustained real-world operation in high stakes scientific environments.
III.1 Task Extraction: From Conversation to Actionable Tasks
LM-driven assistants often operate over extended, multi-turn conversations. While these transcripts capture the full context of an interaction, they are not directly actionable; unfiltered histories can clutter downstream prompts, inflate token usage, and obscure the key objectives. Our framework addresses this challenge through LM-powered transformation of conversational context into structured, machine-readable tasks.
The process begins with intelligent context compression, in which the LM analyzes long conversation histories across multiple turns to isolate the relevant information, remove redundancy, and detect implicit requirements embedded in conversational patterns. This ensures that only essential, high-value context is carried forward.
Extraction is further strengthened through multi-source data integration, combining personalized history from user memory systems, domain-specific knowledge bases for terminology resolution, and external resources such as databases, APIs, and file repositories. During this stage, the system also detects inter-task dependencies, enabling coherent sequencing in downstream workflows.
Finally, task formalization converts free-form natural language requests into explicit, well-defined objectives with clear constraints and dependency structures. This structured representation ensures that the downstream components can execute tasks efficiently and reliably.
III.2 Capability Classification: Adaptive Tool Selection
As the number of available tools grows, LM-based systems face an increasing risk of prompt explosion, a condition where the accumulation of tool-related and contextual information overwhelms the model’s input. This challenge is especially pronounced in niche domains, where reliable performance cannot be achieved through general language understanding alone; instead, prompts must include detailed expert knowledge and domain-specific documentation, further amplifying input size and complexity. Our framework addresses this problem through per-capability relevance analysis, ensuring that only tools pertinent to the current task are considered.
The process begins with individual capability analysis, in which each available tool is evaluated independently for task relevance. This evaluation is framed as a binary classification problem (relevant or irrelevant) and leverages few-shot learning with capability-specific examples. Relevance judgments are context-sensitive, drawing on the explicit task descriptions extracted from conversations and using capability-specific examples and instructions for binary classification decisions.
This targeted filtering enables prompt explosion mitigation by passing only the documentation and interface details of classified-relevant capabilities to the orchestrator. As a result, prompt complexity is decoupled from the total number of available capabilities, allowing practical deployment even with extensive capability inventories.
III.3 Execution Planning: Upfront Orchestration
Standard ReAct-style agents invoke tools reactively, reasoning and acting in short cycles. While effective for simple tasks, this approach can become brittle in complex workflows involving multiple capabilities. Without a global view of the process, such agents risk losing track of overarching objectives, calling tools in suboptimal sequences, or overlooking critical dependencies. Errors in early steps can propagate unchecked, redundant tool calls can waste resources, and there are few natural points for human oversight before costly or irreversible actions are taken.
Our framework addresses these issues with a plan-first architecture, generating a complete execution plan before any tool is called. Multi-step workflows are coordinated with explicit input–output dependencies through context-key mapping, and planning intelligence is separated from execution logic. Plans can be serialized for inspection, modification, and resumption, with an optional planning mode that presents the plan to the user for approval or adjustment before execution begins.
III.4 Execution: Production-Ready Reliability
Operating in high-stakes domains requires systems that go beyond basic orchestration, combining resilience to failures with clear operator control and flexible deployment. The framework achieves this by building on LangGraph’s checkpointing for state continuity, while adding structured error classification and recovery strategies. Errors are handled through bounded retries, re-planning, or reclassification, allowing execution to adapt gracefully without sacrificing reliability.
Human-in-the-loop approval is built in as a first-class capability rather than an afterthought. Structured interrupts allow operators to inspect and approve plans, code, or memory operations before side effects occur, with all state preserved for resumption. This ensures that sensitive operations remain auditable and controllable, while still benefiting from automation when oversight is not required.
The execution layer also emphasizes scientific usability. Python services support seamless switching between local and containerized execution, making it straightforward to adapt pipelines from development laptops to HPC clusters. Generated code is automatically packaged into Jupyter notebooks for inspection and editing, while artifact management captures intermediate and final outputs. Together with multiple interface options, command-line tools, browser-based workflows, and containerized deployments, these features make the framework both reliable for production use and practical for scientific environments.
IV Case Studies
We illustrate the Alpha Berkeley Framework through two complementary case studies. The first is a wind farm monitoring example, presented as a illustrative tutorial example to demonstrate how the framework decomposes natural language requests into structured workflows. The second is a real deployment at the ALS, a U.S. Department of Energy synchrotron facility, where the framework is integrated directly into the accelerator control room. Together, these examples highlight both the general applicability of the framework and its readiness for production use in safety-critical scientific environments.
IV.1 Wind Farm Example
The wind turbine monitoring example demonstrates our framework’s capabilities through a multi-domain analysis workflow. The vital part of this example is the framework’s capacity to showcase the automatic orchestration, i.e. picking the right agent, executing the right agent with minimum overhead.
The shown study validates three key aspects of the framework’s approach to workflow orchestration: (1) automatic dependency resolution in multi-step processes, (2) quick and easy integration of domain knowledge with computational analysis, and (3) production-ready deployment patterns with human oversight.
Consider following user request:
Our wind farm has been underperforming lately. Can you analyze the turbine performance over the past 2 weeks, identify which turbines are operating below industry standards, and rank them by efficiency? I need to know which ones require immediate maintenance attention.
This example shows how our framework automatically decompose this query into a coordinated execution plan that transforms this user request into an transparent, interpretable workflow of steps that are using autonomous agents for temporal reasoning, operating with diverse data sources, use domain expertise and even perform an data analysis, all while maintaining transparency and human oversight necessary for reliable industrial applications.
The orchestrator identifies six distinct capabilities, shown in Figure 2, required for completion: temporal parsing (Step 1), source data integration (Steps 2-3), knowledge extraction (Step 4), computational analysis (Step 5), and response generation (Step 6). Notably, the framework recognizes that turbine performance analysis requires both operational data and environmental context (weather conditions) without explicit user instruction.
It demonstrates the framework’s ability to coordinate heterogeneous data sources within a unified workflow. In the workflow, the Steps 2-4 retrieve turbine sensor data, weather measurements, and extract performance benchmarks from technical documentation. The knowledge retrieval step (Step 4) highlights the framework’s ability to integrate external data sources, incorporating databases and knowledge repositories via the data source provider pattern. The knowledge provider retrieves numerical parameters and structured data that are formatted for direct consumption by downstream Python code generation, showcasing integration of external knowledge systems into dynamic analysis workflows.
In order to reliably generate Python code, we do automated hierarchical analysis planning for a given user’s query followed by dynamic code generation. First, language model creates a structured computational plan that breaks the analysis down into discrete phases (data preparation, performance metrics calculation, industry benchmark comparison). This plan is then converted into executable Python code that properly accesses the structured context data from the previous steps.
Task Classification $\rightarrow$ 6 capabilities identified Execution Planning $\rightarrow$ 6-step plan generated Step (1/6) time_range_parsing $\neg$ Input: past two weeks $\neg$ Output: 2025-07-26 to 2025-08-09 Step (2/6) turbine_data_archiver $\neg$ Source: Mock Turbine API, time range $\neg$ Output: 1,680 turbine readings retrieved. Step (3/6) weather_data_retrieval $\neg$ Source: Mock Weather API $\neg$ Output: 336 wind speed measurements retrieved Step (4/6) knowledge_retrieval $\neg$ Source: Knowledge Base (LLM processing) $\neg$ Output: Performance thresholds 100%-85% $\rightarrow$ excellent 75%-85% $\rightarrow$ good 0%-75% $\rightarrow$ maintenance Step (5/6) turbine_analysis $\neg$ Process: LLM creates analysis plan $\neg$ Execute: Python code generation & execution $\neg$ Output: Results calculated and ranked Step (6/6) respond $\neg$ Input: Analysis results + knowledge threshold $\neg$ Output: Maintenance report with turbine rankings
Figure 2: Overview of the task classification and execution planning process. The diagram illustrates the automatically generated sequential steps involved in analyzing and execution of tasks. Each step is associated with specific inputs and outputs, demonstrating the flow of information and decision-making in the system.
The generated code demonstrates the framework’s type-safe context propagation system, where the code generator derives correct data access patterns directly from context type definitions, for example, pd.DataFrame({’timestamp’: context.TURBINE_DATA.key.timestamps}). This approach eliminates a major source of runtime errors common in traditional tool-calling frameworks and avoids the pitfalls of manual data serialization.
Furthermore, before executing generated code, the framework integrates human approval, which is critical, particularly for sensitive operations.
IV.2 Advanced Light Source Deployment
The ALS at Lawrence Berkeley National Laboratory is a U.S. Department of Energy synchrotron facility that has provided bright X-ray and ultraviolet light to a broad scientific community for more than three decades [42]. In this case study, the Alpha Berkeley Framework is applied in the ALS control room, where the accelerator is operated and optimized to deliver beams to over 40 experimental stations. This environment poses distinctive challenges: real-time integration with control systems, coordination across many heterogeneous subsystems, and strict adherence to safety protocols.
Consider the following user request:
"Get the minimum and maximum value of all ID gap values in the three days. Then write a script which moves each ID from maximum to minimum gap and back while measuring the vertical beam size at beamline 3.1. Sample the gap range with 30 points, wait 5s after each new setpoint for the ID to settle and measure the beamsize 5 times at 5Hz. Return a hysteresis plot beam size vs gap."
This query illustrates the kind of complex, multi-step orchestration that the framework supports. It requires historical data analysis, parameter sweeps, synchronized diagnostics, and scientific visualization, all subject to operational safety requirements in a production environment. Tasks of this kind would typically require an expert operator a couple of hours to identify the correct control system variables, retrieve historical data, and handcraft the measurement scripts. By contrast, the framework generates and executes the entire workflow automatically from a single natural language request.
A distinctive aspect of this example is the chaining of multiple Python-based capabilities generated dynamically by the system. Historical data analysis is carried out by Python code produced on the fly by a language model, with its results passed into a subsequent Python generation step that creates the measurement script. The script embeds these values into a safe control procedure. Crucially, this process is orchestrated in a general way rather than relying on fixed templates, demonstrating that the framework can reliably compose complex, multi-stage Python workflows in unpredictable scenarios.
The deployment integrates a diverse set of data and control sources. The system resolves user instructions into relevant control system variables out of a space of more than 10,000 channels, retrieves historical time-series data spanning years of operations, and interacts in real time with the accelerator’s process control layer. Each of these services is encapsulated as a capability that can be invoked, combined, and reused within an execution plan.
Safety-critical operation management is embedded throughout. All machine control steps require explicit human approval, and planning mode presents execution plans for review and modification before execution. Safe operating boundaries prevent unauthorized changes, while containerized execution environments enforce strict access control and isolate machine-interfacing code.
In the hysteresis analysis, task classification identified the necessary capabilities: parameter discovery, historical data retrieval, data analysis, machine operation, and visualization. The orchestrator produced a structured plan that combined semantic mapping of control system variables, range determination from historical archives, dynamic generation of a safe measurement script, and automated visualization. Each step produced structured context consumed by the next, enabling the framework to transform a high-level request into a multi-stage experimental procedure with traceable intermediate artifacts.
Figure 3 shows that the beam size remains stable across the full range of insertion device gaps, which is the correct physical outcome. The framework carries out the entire experiment automatically: it finds the right control variables, pulls historical data to define the scan range, generates a safe measurement script, runs the scan, and produces an annotated plot. From a single natural-language request, the system completes a multi-step procedure that normally requires an expert hours to set up.
Figure 3: Hysteresis measurement of insertion device gap versus vertical beam size at the ALS. The execution plan generated by the Alpha Berkeley Framework combined historical range extraction, automated script generation, and real-time machine control. The agent performed a 30-point bidirectional gap sweep with 5 repeated measurements per point, producing the plot shown here for one device.
V Conclusion
We have presented the Alpha Berkeley Framework, a distributed multi-agent architecture designed for complex, information-driven workflows in heterogeneous scientific environments. By dynamically assembling planning agents from capability-specific modules, the framework achieves both modularity and scalability. It supports the generation and sharing of intermediate artifacts, such as code and structured data, that can be consumed by other agents, enabling flexible orchestration and emergent workflows. Human oversight is embedded as a first-class feature, with planning mode and structured interrupts ensuring safety and transparency in high-stakes operations.
The framework has been successfully demonstrated at the ALS, where it executed a multi-step analysis and control procedure directly on the accelerator. This deployment illustrates how agentic systems can enhance the usability, reliability, and efficiency of facility operations by translating natural language instructions into reproducible experimental workflows.
Looking ahead, the framework is being applied beyond accelerator operations to beamlines at the ALS, where it is used to integrate data acquisition, analysis, and control. These ongoing efforts highlight its readiness for routine deployment in production environments and its versatility as a facility-wide capability for scientific user environments.
Acknowledgments
This research used the CBorg AI platform and resources provided by the IT Division at Lawrence Berkeley National Laboratory. We gratefully acknowledge Andrew Schmeder for his consistent responsiveness and support, which ensured that CBorg served as an invaluable resource for the development of this framework and the ALS expert agent.
We are grateful to Alex Hexemer for his steady encouragement and advocacy in advancing agentic AI as a facility-wide capability at the ALS, which strongly influenced the broader direction of this framework.
We also thank Frank Mayet for his collaborative spirit and for sharing insights from his pioneering Gaia prototype, which helped guide the early trajectory of agentic AI efforts at the ALS.
We warmly thank our colleagues at the ALS, Fernando Sannibale, Marco Venturini, Simon Leemann, Drew Bertwistle, Erik Wallen, Hiroshi Nishimura, Tom Scarvie, Christoph Steier, Tynan Ford, and Edison Lam, for their expertise, feedback, and support in exploring and shaping the integration of the framework into accelerator operations.
We further acknowledge the use of AI tools during the preparation of this work. Cursor, primarily with Claude 4, was employed extensively during development, while GPT-5 was used to refine the language of this manuscript.
This work was supported by the Director of the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.